### v1

# On Mixed-Precision Iterative Methods and Analysis for Nearly Completely Decomposable Markov Processes

Abstract
In this paper we consider the problem of computing the stationary distribution of nearly completely decomposable Markov processes, a well-established area in the classical theory of Markov processes with broad applications in the design, modeling, analysis and optimization of computer systems. We design general classes of algorithmic solution approaches that exploit forms of mixed-precision computation to significantly reduce computation times and that exploit forms of iterative approximate methods to mitigate the impact of inaccurate computations, further reduce computation times, and ensure convergence. Then we derive a mathematical analysis of our general algorithmic approaches that establishes theoretical results on approximation errors, convergence behaviors, and other algorithmic properties. Numerical experiments demonstrate that our general algorithmic approaches provide significant improvements in computation times over the most efficient existing numerical methods.
1 Introduction
Markov processes often play an important role in the design and mathematical modeling, analysis and optimization of a wide range of computer systems. A critically important goal in the mathematical modeling, analysis and optimization of such computer systems and their dynamic behavior concerns the need to determine the stationary distribution of the corresponding Markov process, from which various measures and quantities of interest can be directly obtained to facilitate and support the design, analysis and optimization of computer systems and applications.
More precisely, consider a discrete-time Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ defined on the state space $[n]:=\{1,\ldots,n\}$, and let ${\mathbf{P}}\in\mathbb{R}^{n\times n}$ and $\bm{\pi}:=(\pi_{1},\ldots,\pi_{n})$ respectively denote its transition probability matrix and stationary distribution. Here, $\mathbb{Z}_{+}$ and $\mathbb{R}$ denote the set of nonnegative integers and real numbers, respectively. Assuming this Markov process is irreducible and ergodic, then the invariant probability vector $\bm{\pi}$ exists and is uniquely determined as the solution of [25, 26]
$\bm{\pi}=\bm{\pi}\mathbf{P}\qquad\mbox{and}\qquad\bm{\pi}\mathbf{1}=\|\bm{\pi}\|_{1}=1,$ (1)
[…]
One particularly important class of structural properties often arising in such Markov processes concerns the so-called “aggregation of variables” [22] in which all the states (variables) of the system can be grouped into a relatively small number of relatively large clusters such that the magnitude of transitions between states within a cluster overwhelmingly dominate the relatively tiny magnitude of transitions between states in different clusters. Markov processes with these structural properties, called nearly completely decomposable (NCD), make it possible to effectively investigate the system and its dynamics by exploiting the analysis of interactions among the states within each individual cluster without reference to the interactions among the states across clusters and by exploiting the analysis of interactions among the states across different clusters without reference to the interactions among the states within clusters. Such NCD properties arise naturally in many computer systems and applications due to the prevalence of levels of modular abstraction or hierarchies of components in the design of these computer systems and applications where interactions within modules or components dominate interactions between modules or components. In addition to the various traditional Markov process instances within computer systems (see, e.g., [8, 10, 9, 6]), these NCD structural properties arise naturally in various traditional aspects of many other domains such as biology, control theory, dynamical systems, economics, genetics, inventory and production systems, queueing theory, and social sciences; refer to, e.g., [5, 22, 11, 4, 9, 6, 3, 30].
More formally, consider a discrete-time Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ defined on the state space $[n]$ with transition probability matrix $\mathbf{P}$ taking the block structural form
$\mathbf{P}=\begin{pmatrix}\mathbf{P}_{11}&\mathbf{P}_{12}&\cdots&\mathbf{P}_{1m}\\ \mathbf{P}_{21}&\mathbf{P}_{22}&\cdots&\mathbf{P}_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ \mathbf{P}_{m1}&\mathbf{P}_{m2}&\cdots&\mathbf{P}_{mm}\end{pmatrix},$ (2)
where the $\mathbf{P}_{ij}$, $i,j\in[m]$, are nonnegative block matrices in $\mathbb{R}^{n_{i}\times n_{j}}$ such that $\sum_{j\in[m]}\mathbf{P}_{ij}$ are stochastic (i.e., rows sum to one with each element in $[0,1]$), and $\mathbf{P}$ is a stochastic matrix in $\mathbb{R}^{n\times n}$, $n=\sum_{i\in[m]}n_{i}$. Define $\sigma_{i}:=\sum_{\ell=1}^{i-1}n_{\ell}$, $i\in[m]$. Under the assumption that the Markov process is NCD, the elements of the off-diagonal block matrices $\mathbf{P}_{ij}$ are very small relative to the elements of the diagonal block matrices $\mathbf{P}_{ii}$, $i,j\in[m],i\neq j$. In particular, we assume throughout that
$\|\mathbf{P}_{ii}\|=O(1),\qquad\|\mathbf{P}_{ij}\|=O(\epsilon),\qquad\qquad i,j\in[m],\;i\neq j,$ (3)
[…]
$\bm{\pi}:=(\bm{\pi}_{1},\bm{\pi}_{2},\ldots,\bm{\pi}_{m}),\qquad\bm{\pi}_{i}:=(\pi_{i1},\pi_{i2},\ldots,\pi_{in_{i}}),\qquad\pi_{ik}:=\lim_{s\rightarrow\infty}\mathbb{P}[X(s)=(\sigma_{i}+k)],$ (4)
for $k\in[n_{i}]$, $i\in[m]$. The limiting probability vector $\bm{\pi}$ is the stationary distribution of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$. We assume throughout that this process is irreducible and ergodic, and thus its invariant probability vector $\bm{\pi}$ exists and is uniquely determined as the solution of (1).
In addition to their broad applications across a wide range of domains in computing, engineering, science and business, NCD structural properties represent an important fundamental area in the classical theory of Markov processes [22, 27, 28, 18, 19, 7, 14, 24, 30]. Hence, several different (yet related) numerical methods have been developed to efficiently compute the stationary distribution $\bm{\pi}$ of NCD Markov processes, with the most efficient based on different forms of exploiting the aggregation and disaggregation of NCD structural properties as part of an iterative computational method. This collection of the most efficient numerical methods for NCD Markov processes consists predominantly of the algorithms due to Takahashi [27], Vantilborgh [28], and Koury, McAllister and Stewart [18]. On the one hand, although they differ in terms of specific details, these methods exploit aggregation-disaggregation as part of an iterative computational process in a very similar manner and with similar convergence behaviors [7]. On the other hand, despite their significant performance benefits over directly solving (1), these numerical methods can still be prohibitively expensive for computing the stationary distribution of highly large-scale NCD Markov processes and for supporting online or real-time applications of such stochastic processes. Moreover, all of these most efficient methods involve computations in full precision and none of them take advantage of lower precision technology.
Our Contributions. Our focus in this paper is on the design and analysis of numerical methods for computing the stationary distribution $\bm{\pi}$ of NCD Markov processes that address the performance bottlenecks of existing approaches and provide significant computational improvements to support the very large scale of current and emerging computer systems and applications as well as their potential deployment in online or real-time environments. In doing so, we make the following important technical contributions. We first design general classes of computational algorithmic approaches to efficiently determine the solution of (1) for the invariant probability vector $\bm{\pi}$ in (4) of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ whose $\mathbf{P}$ matrix has the structural form (2) with properties (3). Each of our general approaches involve a combination of different forms of exploiting advances in computer architectures related to mixed-precision computation, which significantly reduce computation times at the expense of inaccuracies, and exploiting iterative approximate computing methods, which mitigate the impact of inaccurate computations, further reduce computation times, and guarantee convergence. It is important to note that, although low- and mixed-precision computation have received considerable attention in machine learning and related applications, such approaches and methods are far less well established in the context of numerical solutions of stochastic processes. We then derive a mathematical analysis of our general classes of algorithmic approaches that rigorously establishes their important theoretical properties including results on algorithmic approximation errors and convergence behaviors. Lastly, we conduct numerous numerical experiments which empirically demonstrate that our general algorithmic approaches provide orders of magnitude improvements in the computation times to obtain the stationary distribution of NCD Markov processes over the most efficient existing methods which are based on aggregation and disaggregation.
Meanwhile, we note that Horton and Leutenegger [15], motivated by multigrid methods, propose an algorithmic approach for computing the stationary distribution of Markov processes in general without any special structural properties. Although the proposed algorithm is somewhat related to aggregation-disaggregation methods, with the latter methods seen as a special case under specific algorithmic choices, Horton and Leutenegger do not provide any theoretical results on convergence or other algorithmic properties, and they only provide empirical results for Markov processes without NCD structural properties [15]. For these reasons, we do not consider further in this paper the multigrid-based method of Horton and Leutenegger. It is important to note, however, that our general algorithmic approaches devised in this paper within the context of aggregation-disaggregation methods can be analogously exploited within the context of their multigrid-based method for more general Markov processes beyond those having NCD structural properties.
Organization of paper. Section 2 provides technical preliminaries on NCD Markov processes, followed by the presentation of our design of algorithmic solution approaches in Section 3. We derive our mathematical analysis of these algorithmic approaches in Section 4, followed by our proofs of the corresponding theoretical results in Section 5. The empirical results from a representative sample of numerical experiments are presented in Section 6, followed by concluding remarks.
2 Technical Preliminaries
In this section, we present preliminary technical details and results for NCD Markov processes, including additional standard conditions and properties assumed throughout the paper. We shall use the following standard notation. Let $\mathbb{R}$, $\mathbb{Z}_{+}$ and $\mathbb{N}$ denote the set of real numbers, nonnegative integers and natural numbers, respectively. Define $[n]:=\{1,\ldots,n\}=\mathbb{N}\setminus\{n+1,n+2,\ldots\}$. Bold lowercase variables denote vectors, while bold uppercase variables denote matrices. Let $\mathbf{I}$ denote the identity matrix of appropriate dimension with $\mathbf{I}_{n}$ denoting the $n\times n$ identity matrix.
Beyond our assumptions that the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ has the structural form (2) with properties (3) and is irreducible and ergodic, we further assume that $z\leq\|\bm{\pi}_{i}\|_{1}$ for a constant $z>0$, and thus from (1) we have $10<z\leq\|\bm{\pi}_{i}\|_{1}<$ 1, $i\in[m]$. The transition probability matrix $\mathbf{P}$, since it is a stochastic matrix, has a dominant unit eigenvalue and corresponding right eigenvector $\mathbf{1}$ given by
$\mathbf{P}\mathbf{1}=\mathbf{1}.$
We additionally assume no other eigenvalues of $\mathbf{P}$ have absolute value one.
Turning to the properties of the eigenvalues of the diagonal block matrices, let $\mathbf{v}_{i}>0$ denote the left eigenvector of $\mathbf{P}_{ii}$ that corresponds to its dominant eigenvalue $\lambda_{i_{1}}$, $i\in[m]$. We assume that
$\lambda_{i_{1}}=1-{g}_{i}\epsilon+o(\epsilon),\qquad{g}_{i}>0,\;\;i\in[m],$
i.e., $\mathbf{P}_{ii}$ has only one dominant eigenvalue that is close to one. We further assume that the remaining eigenvalues of $\mathbf{P}_{ii}$ are uniformly bounded away from one. Then, it is well known [19] that there exists a nonsingular matrix $\mathbf{W}_{i}=\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}$ such that
$\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}\mathbf{P}_{ii}=\begin{pmatrix}1-{g}_{i}\epsilon+o(\epsilon)&0\\ \mathbf{0}&\mathbf{T}_{i}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix},$ (5)
with the norms of $\mathbf{W}_{i}$ and $\mathbf{W}_{i}^{-1}$ bounded by a constant independent of $\epsilon$, $i\in[m]$. From (5), we derive
$\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})$ $=\begin{pmatrix}{g}_{i}\epsilon+o(\epsilon)&0\\ \mathbf{0}&\mathbf{I}-\mathbf{T}_{i}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}\qquad\mbox{ and}$
$\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})^{-1}$ $=\begin{pmatrix}({g}_{i}\epsilon+o(\epsilon))^{-1}&0\\ \mathbf{0}&(\mathbf{I}-\mathbf{T}_{i})^{-1}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}.$ (6)
By further imposing the condition
$\|(\mathbf{I}-\mathbf{T}_{i})^{-1}\|\leq{w}_{i},$ (7)
for a constant ${w}_{i}$ independent of $\epsilon$, we guarantee the nondominant eigenvalues $\lambda_{i_{2}},\ldots,\lambda_{i_{n_{i}}}$ of $\mathbf{P}_{ii}$ are uniformly bounded away from one, $i\in[m]$.
Now, define
$\hat{\bm{\pi}}:=(\hat{\bm{\pi}}_{1},\hat{\bm{\pi}}_{2},\ldots,\hat{\bm{\pi}}_{m}),\qquad\hat{\bm{\pi}}_{i}:=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}},\qquad\mathbf{R}_{ij}:=\hat{\bm{\pi}}_{i}\mathbf{P}_{ij}\mathbf{1},$ (8)
for $i,j\in[m]$. It follows from our assumptions on the matrix $\mathbf{P}$ and from $\|\hat{\bm{\pi}}_{i}\|_{1}=1$, $i\in[m]$, that $\mathbf{R}$ is an irreducible stochastic matrix with unique left eigenvector $\mathbf{s}>0$ corresponding to the unit eigenvalue. More precisely, we have
$\mathbf{s}=\mathbf{s}\mathbf{R}\qquad\mbox{and}\qquad\mathbf{s}\mathbf{1}=\|\mathbf{s}\|_{1}=1.$ (9)
Further observe that
$\|\bm{\pi}_{j}\|_{1}=\bm{\pi}_{j}\mathbf{1}=\sum_{i=1}^{m}\bm{\pi}_{i}\mathbf{P}_{ij}\mathbf{1}=\sum_{i=1}^{m}\|\bm{\pi}_{i}\|_{1}\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}\mathbf{P}_{ij}\mathbf{1}=\sum_{i=1}^{m}\|\bm{\pi}_{i}\|_{1}\mathbf{R}_{ij}$
together with $\sum_{i=1}^{m}\|\bm{\pi}_{i}\|_{1}=1$, and thus the vector $(\|\bm{\pi}_{1}\|_{1},\|\bm{\pi}_{2}\|_{1},\ldots,\|\bm{\pi}_{m}\|_{1})$ satisfies (9), which as a result of the uniqueness of its solution yields
$\mathbf{s}=(\|\bm{\pi}_{1}\|_{1},\|\bm{\pi}_{2}\|_{1},\ldots,\|\bm{\pi}_{m}\|_{1}).$ (10)
Next, it follows from $\mathbf{P}$ being an irreducible matrix that there exist uniformly bounded matrices $\mathbf{X}$ and $\mathbf{Y}$ such that
$\mathbf{S}:=(\mathbf{1},\mathbf{X})=\begin{pmatrix}\bm{\pi}\\ \mathbf{Y}\end{pmatrix}^{-1},\qquad\mathbf{S}^{-1}\mathbf{R}\mathbf{S}=\begin{pmatrix}1&0\\ 0&\mathbf{B}\end{pmatrix},\qquad\|\mathbf{S}\|=O(1)=\|\mathbf{S}^{-1}\|.$ (11)
By further assuming
$\|(\mathbf{I}-\mathbf{B})^{-1}\|=O\left(\frac{1}{\epsilon}\right),$ (12)
we guarantee the nondominant eigenvalues of the aggregation matrix $\mathbf{R}$ are bounded away from one by an amount proportional to $\epsilon$.
3 Algorithmic Solutions
In this section, we consider various algorithmic approaches for computing the solution of (1) to determine the stationary distribution $\bm{\pi}$ of NCD Markov processes with the goal of providing significant improvements in computational and theoretical properties over the most efficient numerical methods for doing so. Both of our main algorithmic approaches are based on modified variants of the algorithm due to Koury, McAllister and Stewart [18], as the most recent of the three primary aggregation-disaggregation numerical methods, in which we sharply reduce the computational bottlenecks associated with solving systems of linear equations by exploiting various general classes of iterative approximate mixed-precision computing methods. The primary differences between our two main algorithmic approaches concern the aggressiveness of the degree to which we exploit the various general classes of iterative approximate mixed-precision computing methods.
Since both of our algorithmic approaches are based on modified variants of the Koury-McAllister-Stewart (KMS) algorithm, we first present the main steps the original algorithm. The KMS algorithm is composed of an iterative numerical method that starts at iteration 00 with any given initial approximation $\bm{\pi}^{(0)}=(\bm{\pi}_{1}^{(0)},\ldots,\bm{\pi}_{m}^{(0)})$ of the desired solution $\bm{\pi}$. Then, for each subsequent iteration $t=1,2,3,\ldots$, Step $\mathbf{1}$ consists of normalizing the vector components of the current estimate $\bm{\pi}^{(t-1)}$ of the solution $\bm{\pi}$, i.e., compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\bm{\pi}_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1}$ according to the middle equation in (8), $i\in[m]$. Step $\mathbf{2}$ comprises obtaining the elements of the aggregation matrix $\mathbf{R}^{(t-1)}$ leading up to the current iteration, i.e., compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$ according to the rightmost equation in (8), $i,j\in[m]$. Next, Step $\mathbf{3}$ consists of obtaining the dominant left eigenvector $\mathbf{s}$ of $\mathbf{R}$ by computing the solution of $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ according to (9). Step $\mathbf{4}$ comprises computing the Hadamard product $\mathbf{z}^{(t)}=\mathbf{s}^{(t-1)}\odot\hat{\bm{\pi}}^{(t-1)}=(s_{1}^{(t-1)}\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,s_{m}^{(t-1)}\hat{\bm{\pi}}_{m}^{(t-1)})$. Then, Step $\mathbf{5}$ consists of solving the sequence of $m$ systems of linear equations $\bm{\pi}_{i}^{(t)}=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j<i}\mathbf{z}_{j}^{(t)}\mathbf{P}_{ji}+\sum_{j>i}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$ in the order $i=m,\ldots,1$, thus rendering the blockwise estimates of the components of $\bm{\pi}^{(t)}$ for iteration $t$. Step $\mathbf{6}$ comprises conducting a test for convergence, and halting with the return of the current estimate $\bm{\pi}^{(t)}$ if the solution is sufficiently accurate; otherwise, the iteration index $t$ is incremented and the iterative process of Step $\mathbf{1}$ – Step $\mathbf{6}$ is repeated.
We now present each of our two main algorithmic approaches in turn.
3.1 Conservatively Modified Koury-McAllister-Stewart Algorithm
The performance bottlenecks of the KMS algorithm concern computing the solution of $m+1$ systems of linear equations in each iteration, specifically the sequence of $n_{i}\times n_{i}$ linear systems in Step $\mathbf{5}$, $i\in[m]$, as well the $m\times m$ linear system in Step $\mathbf{3}$ for problems with large $m$. To address these performance bottlenecks, our first general class of modified variants of the KMS algorithm consists of replacing the standard linear solvers in Step $\mathbf{5}$ and Step $\mathbf{3}$ with general classes of approximate mixed-precision computing methods, each of which involves a combination of different forms of lower precision technology to reduce computation costs and different forms of iterative approximate methods to further reduce computation costs with adjustments to realize full-precision results. Instances of the former include forms of multi-precision arithmetic or stochastic rounding, whereas instances of the latter include forms of iterative refinement [29], Richardson iteration [13], and generalized minimal residual (GMRES) [21] methods.
Our first general class of algorithmic approaches based on modified variants of the KMS algorithm, as summarized in Algorithm 1, can be applied with any instances of the general classes of approximate mixed-precision computing methods to solve the linear systems in Step $\mathbf{5}$ and Step $\mathbf{3}$. In particular, we replace the single invocation of a full-precision linear system solver in the original KMS algorithm with an appropriate iterative approximate mixed-precision linear system solver of interest involving a sequence of steps that lead to full-precision accuracy. To reduce the computation costs as much as possible while still realizing full-precision results, we determine a level of reduced precision (such as single-precision, half-precision, bfloat16, and so on) based on the properties of the linear system that ensures convergence under the approximate mixed-precision computing method of interest to yield full-precision accuracy. As a representative example, we shall focus on the iterative refinement method in which case the selection of the proper reduced precision is based on an estimate of the condition number and the norm of the matrix $\mathbf{P}_{ii}$ for each linear system in Step $\mathbf{5}$, $i\in[m]$, and similarly for the linear system of Step $\mathbf{3}$. Specifically, for each linear system with matrix $\mathbf{P}_{ii}$, we determine the machine precision $\delta$ such that $\delta\kappa(\mathbf{P}_{ii})\|\mathbf{P}_{ii}\|$ is sufficiently small to ensure convergence [20], where $\kappa(\mathbf{P}_{ii})$ denotes the condition number of the matrix $\mathbf{P}_{ii}$, $i\in[m]$.
Once the level of reduced precision has been determined in this manner, the iterative refinement method is used to solve each of the linear systems in Step $\mathbf{5}$ and Step $\mathbf{3}$, generically denoted by $\mathbf{A}\mathbf{x}=\mathbf{b}$. At every step $k$ of the iterative refinement method within each outer-loop iteration $t$ of Algorithm 1, we compute the residual $\mathbf{d}_{k}=\mathbf{b}-\mathbf{A}\mathbf{x}_{k}$ and then solve for $\mathbf{y}_{k}$ in $\mathbf{A}\mathbf{y}_{k}=\mathbf{d}_{k}$ using the reduced precision method. The resulting solution $\mathbf{y}_{k}$ is added to the previous estimate $\mathbf{x}_{k}$ to obtain $\mathbf{x}_{k+1}=\mathbf{x}_{k}+\mathbf{y}_{k}$, and the step index $k$ is incremented. This process of inner-loop steps is repeated until the residual falls below the error of double-precision arithmetic, which is guaranteed by our selection of the reduced precision level [20]. An analogous approach can be followed for any of the approximate mixed-precision computing method alternatives to iterative refinement.
Algorithm 1 Conservatively Modified KMS Algorithm
Input: Irreducible stochastic matrix $\mathbf{P}$
Output: Estimate of probability distribution $\bm{\pi}$ such that $\bm{\pi}=\bm{\pi}\mathbf{P}$ and $\bm{\pi}\mathbf{1}=1$
Set $t=1$ and Let $\bm{\pi}^{(0)}$ be the given initial approximation to the solution $\bm{\pi}$
While: Stopping criteria is not satisfied
 Compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\frac{\bm{\pi}_{i}^{(t-1)}}{\|\bm{\pi}_{i}^{(t-1)}\|_{1}}$, $i\in[m]$, and Obtain $\hat{\bm{\pi}}^{(t-1)}=(\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,\hat{\bm{\pi}}_{m}^{(t-1)})$
 Compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$, $i,j\in[m]$, and Obtain $\mathbf{R}^{(t-1)}$
 Compute solution of linear system $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ using           approximate mixed-precision computing method of interest
 Compute Hadamard product $\mathbf{z}^{(t)}=\mathbf{s}^{(t-1)}\odot\hat{\bm{\pi}}^{(t-1)}$
 Compute solutions of linear systems $\bm{\pi}_{i}^{(t)}=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j<i}\mathbf{z}_{j}^{(t)}\mathbf{P}_{ji}+\sum_{j>i}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$ for $i=m,\ldots,1$           using approximate mixed-precision computing method of interest, and Obtain $\bm{\pi}^{(t)}$
 Conduct convergence test, Setting $t=t+1$ if stopping criteria not satisfied
Return: Estimate $\bm{\pi}^{(t)}$
3.2 Aggressively Modified Koury-McAllister-Stewart Algorithm
Building on the above conservatively modified KMS algorithm to address the performance bottlenecks of the original KMS algorithm, our second general class of algorithmic approaches based on modified variants of the KMS algorithm, as summarized in Algorithm 2, consists of taking a more aggressive approach in exploiting approximate mixed-precision computing methods to replace the standard linear solvers in Step $\mathbf{5}$ and Step $\mathbf{3}$. In particular, Step $\mathbf{5}$ and Step $\mathbf{3}$ can be computed with fewer iterations at the expense of not realizing full-precision results which is then addressed through appropriated levels of preconditioning and iteration for further adjustments. We shall solely focus on Step $\mathbf{5}$ in this section with the understanding that the same approximate mixed-precision computing approach can be similarly applied in Step $\mathbf{3}$;. alternatively, Step $\mathbf{3}$ can be computed by employing our approximate mixed-precision computing approach in Algorithm 1.
Define the block matrices $\mathbf{D}_{ii}=\mathbf{I}_{n_{i}}-\mathbf{P}_{ii}$, $i\in[m]$; the strictly block-lower-triangular matrices $\mathbf{L}_{ij}=\mathbf{P}_{ij}$, $i>j$, and $\mathbf{L}_{ij}=\mathbf{0}$, $i\leq j$, $i,j\in[m]$; and the strictly block-upper-triangular matrices $\mathbf{U}_{ij}=\mathbf{P}_{ij}$, $i<j$, and $\mathbf{U}_{ij}=\mathbf{0}$, $i\geq j$, $i,j\in[m]$. Further define the block-diagonal matrix $\mathbf{D}=\operatorname*{diag}(\mathbf{D}_{11},\ldots,\mathbf{D}_{mm})$. We then can decompose the matrix $\mathbf{I}-\mathbf{P}$ as
$\mathbf{I}-\mathbf{P}=\mathbf{D}-\mathbf{L}-\mathbf{U}.$
Now, define the diagonal matrix $\mathbf{D}^{(t-1)}$ of iteration $t-1$ whose corresponding $i$-th principal diagonal block $\mathbf{D}^{(t-1)}_{ii}$ is set to be $\mathbf{D}^{(t-1)}_{ii}=(s_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1})\,\mathbf{I}_{n_{i}}$, $i\in[m]$. From Step $\mathbf{4}$, we then can write $\mathbf{z}^{(t)}=\bm{\pi}^{(t-1)}\mathbf{D}^{(t-1)}$ which, in combination with the systems of linear equations in Step $\mathbf{5}$ of the original KMS algorithm, leads to $\bm{\pi}^{(t)}=\mathbf{z}^{(t)}\mathbf{L}(\mathbf{D}-\mathbf{U})^{-1}$. Upon substituting the expression for $\mathbf{z}^{(t)}$, we finally have
$\bm{\pi}^{(t)}=\bm{\pi}^{(t-1)}\mathbf{D}^{(t-1)}\mathbf{L}(\mathbf{D}-\mathbf{U})^{-1}.$ (13)
With this as a starting point, our second general class of modified variants of the KMS algorithm is based on exploiting a combination of (13) together with levels of mixed-precision preconditioning and iteration for Step $\mathbf{5}$. In particular, transposing both sides of (13) yields
$\bm{\pi}^{(t)\top}=(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top},$ (14)
from which it is apparent that $\bm{\pi}^{(t)\top}$ can be recast as the solution of a sparse linear system with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as its iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as its right-hand side. Then the main idea of our aggressive mixed-precision computing method for Step $\mathbf{5}$ is to approximately solve the linear systems in (14) via a fixed number of steps of the mixed-precision Richardson iteration method [13]. In essence, the preconditioned Richardson iteration updates the $k_{t}$-th approximation of $\bm{\pi}^{(t)\top}$ via the fixed-point iteration
$\bm{\pi}^{(t)\top}_{k_{t}+1}=(\mathbf{I}-\mathbf{M}^{-1}(\mathbf{D}^{\top}-\mathbf{U}^{\top}))\bm{\pi}^{(t)\top}_{k_{t}}+\mathbf{M}^{-1}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top},$ (15)
where $k_{t}$ denotes the number of mixed-precision Richardson iteration steps in the $t$-th iteration of Algorithm 2 and the matrix $\mathbf{M}^{-1}$ denotes the preconditioner of the iterative method. For example, if $\mathbf{M}\equiv\mathbf{D}^{\top}$, i.e., block-Jacobi preconditioning, the iteration matrix in (15) becomes equal to $\mathbf{D}^{-\top}\mathbf{U}^{\top}$. Throughout the rest of this paper we consider $\mathbf{M}$ to be the product of the LU factors of $\mathbf{D}^{\top}$ obtained using reduced precision. [^1] While not explored in this paper, it is worth mentioning that (15) is equivalent to a linear system with the same iteration matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ and multiple right-hand sides, and thus even further improvements might be possible within the context of our approach; see, e.g., [16, 17].
Algorithm 2 Aggressively Modified KMS Algorithm
Input: Irreducible stochastic matrix $\mathbf{P}$
Output: Estimate of probability distribution $\bm{\pi}$ such that $\bm{\pi}=\bm{\pi}\mathbf{P}$ and $\bm{\pi}\mathbf{1}=1$
Set $t=1$ and Let $\bm{\pi}^{(0)}$ be the given initial approximation to the solution $\bm{\pi}$
Set $\mathbf{I}-\mathbf{P}=\mathbf{D}-\mathbf{L}-\mathbf{U}$ and matrices $\mathbf{L}_{ij}$ and $\mathbf{U}_{ij}$
While: Stopping criteria is not satisfied
 Compute $\hat{\bm{\pi}}_{i}^{(t-1)}=\frac{\bm{\pi}_{i}^{(t-1)}}{\|\bm{\pi}_{i}^{(t-1)}\|_{1}}$, $i\in[m]$, and Obtain $\hat{\bm{\pi}}^{(t-1)}=(\hat{\bm{\pi}}_{1}^{(t-1)},\ldots,\hat{\bm{\pi}}_{m}^{(t-1)})$
 Compute $\mathbf{R}_{ij}^{(t-1)}=\hat{\bm{\pi}}_{i}^{(t-1)}\mathbf{P}_{ij}\mathbf{1}$, $i,j\in[m]$, and Obtain $\mathbf{R}^{(t-1)}$
 Compute solution of linear system $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}^{(t-1)}\mathbf{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ using           approximate mixed-precision computing method of interest
 Compute $\mathbf{D}^{(t-1)}$ whose corresponding $i$-th principal diagonal block $\mathbf{D}^{(t-1)}_{ii}$ is set to be $\mathbf{D}^{(t-1)}_{ii}=$           $(s_{i}^{(t-1)}/\|\bm{\pi}_{i}^{(t-1)}\|_{1})\,\mathbf{I}_{n_{i}}$, $i\in[m]$.
 Compute solution of sparse linear system with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as its iteration matrix and          $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as its right-hand side using approximate mixed-precision computing           method of Richardson iteration (15), and Obtain $\bm{\pi}^{(t)}$
 Conduct convergence test, Setting $t=t+1$ if stopping criteria not satisfied
Return: Estimate $\bm{\pi}^{(t)}$
4 Mathematical Analysis
We now turn to derive a mathematical analysis of the general classes of computational algorithmic approaches of the previous section. Our derivation of an error and convergence analysis is presented first for our conservatively modified KMS algorithmic approach and then for our aggressively modified KMS algorithmic approach, establishing the guaranteed convergence of both of our approaches. We conclude this section with a performance analysis of certain aspects of the computational improvements provided by our exploitation of multi-precision arithmetic in both approaches. The conditions and properties expressed in both Section 2 and the introduction are assumed to hold throughout our analysis.
4.1 Error and Convergence Analysis of Conservatively Modified KMS Algorithm
Our main theoretical result is that Algorithm 1 — our conservatively modified KMS algorithm — is guaranteed to converge with an approximation error that decreases by a factor of $O(\epsilon)$ at each iteration. Recall $\bm{\pi}$ to be the exact stationary distribution of the Markov process $\{X(s)\,;\,s\in\mathbb{Z}_{+}\}$ and $\bm{\pi}^{(t)}$ to be the estimate of the solution $\bm{\pi}$ in iteration $t$. Assuming the estimate $\bm{\pi}^{(t-1)}$ of $\bm{\pi}$ has $O(\varphi)$ accuracy, we first present a lemma that shows the estimate $\mathbf{s}^{(t-1)}$ in iteration $t-1$ of the vector $\mathbf{s}$ to also have $O(\varphi)$ accuracy and further shows the estimate $\mathbf{R}^{(t-1)}$ in iteration $t-1$ of the matrix $\mathbf{R}$ to have $O(\varphi\epsilon)$ accuracy.
Lemma 4.1.
Let $\mathbf{s}=\mathbf{s}\mathbf{R}$ and $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ with $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$, and suppose $\bm{\pi}^{(t-1)}$ is an $O(\varphi)$ approximation to $\bm{\pi}$, $t\in\mathbb{N}$. Then, we have
$\mathbf{s}^{(t-1)}-\mathbf{s}=O(\varphi),\qquad t\in\mathbb{N},$
and the matrix $\mathbf{R}^{(t-1)}$ corresponding to $\bm{\pi}^{(t-1)}$ is an $O(\varphi\epsilon)$ approximation of $\mathbf{R}$, where
$\mathbf{R}^{(t-1)}_{ij}=\frac{\bm{\pi}^{(t-1)}_{i}}{\|\bm{\pi}^{(t-1)}_{i}\|_{1}}\mathbf{P}_{ij}\mathbf{1},\qquad\forall i,j\in[m].$
Recall $\mathbf{v}_{i}$ to be the left eigenvector of $\mathbf{P}_{ii}$ corresponding to its dominant eigenvalue $\lambda_{i_{1}}=1-{g}_{i}\epsilon+o(\epsilon)$, $i\in[m]$. We now establish our main results that shows there is an $O(\epsilon)$ norm difference between the exact stationary distribution $\bm{\pi}$ and the corresponding eigenvectors $\mathbf{v}_{i}$ of $\mathbf{P}_{ii}$ multiplied by an $O(1)$ constant and that shows Algorithm 1 is guaranteed to converge with an approximation error of the estimate $\bm{\pi}^{(t)}$ in each iteration $t$ of the exact solution $\bm{\pi}$ decreasing by a factor of $O(\epsilon)$.
Theorem 4.2.
Let $\mathbf{v}_{i}$ be the left eigenvector of $\mathbf{P}_{ii}$ corresponding to the exact invariant probability vector $\bm{\pi}_{i}$, $i\in[m]$. Then, there exists a constant ${b}_{i}=O(1)$ such that
$\|\bm{\pi}_{i}-{b}_{i}\mathbf{v}_{i}\|_{1}=O(\epsilon).$ (16)
Moreover, Algorithm 1 converges with an approximation error in the solution $\bm{\pi}^{(t)}$ at each iteration $t$ that decreases by a factor of $O(\epsilon)$.
The results of Lemma 4.1 and Theorem 4.2 are based on the assumption that the systems of linear equations in Step $\mathbf{5}$ and Step $\mathbf{3}$ are all solved at the exact accuracy of full precision. Based on our selection of the computational precision with respect to the condition number and the norm of the matrix of the linear system as defined in Section 3.1, we next present a lemma that shows our use of the iterative refinement method in Algorithm 1 is guaranteed to converge with the approximation error of the estimate of the exact solution in each iteration decreasing by a factor of $O(\kappa(\mathbf{A}))$, where $\mathbf{A}$ is the matrix of the linear system and $\kappa(\mathbf{A})$ denotes its condition number.
Lemma 4.3.
Consider computing the solution of the linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$ via the mixed-precision iterative refinement method under the precision selection process employed in Step $\mathbf{5}$ and Step $\mathbf{3}$ of Algorithm 1. Then, this iterative refinement method converges linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each iteration.
Convergence and related results similar to those in Lemma 4.3 hold for the alternative general classes of approximate mixed-precision computing methods of interest. We also note that the linear convergence with a decreasing factor of $O(\kappa(A))$ is typically much faster than the convergence of the outer-loop iteration of Algorithm 1 with a decreasing factor of $O(\epsilon)$.
4.2 Error and Convergence Analysis of Aggressively Modified KMS Algorithm
Our main theoretical result is that Algorithm 2 — our aggressively modified KMS algorithm — is guaranteed to converge with an approximation error that decreases with respect to a tradeoff between an increase in the number of iterations $t$ of the outer loop for fewer numbers of Richardson iteration steps $k_{t}$ in Step $\mathbf{5}$ and a decrease in the computational complexity of the Richardson iterative method for smaller $k_{t}$. To consider these theoretical aspects of our second algorithmic approach, we begin by seeking to understand the effects of stopping the mixed-precision Richardson iteration method after a certain number of steps, say $k\in\mathbb{N}$, as formally expressed in the following lemma.
Lemma 4.4.
Consider computing the solution of the linear system $(\mathbf{D}^{\top}-\mathbf{U}^{\top})\mathbf{x}=\mathbf{b}$ via the Richardson iteration method. Then, after $k$ steps, the absolute error satisfies
$\|\mathbf{x}-\mathbf{x}_{[k]}\|\leq\|\mathbf{D}^{-\top}\|\dfrac{\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|^{k+1}}{1-\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|}\|\mathbf{b}\|.$
We next consider an idealized scenario where $\mathbf{D}^{(t-1)}\equiv\mathbf{C}$, $\forall t\in\mathbb{N}$, and $\mathbf{C}$ is a non-singular matrix having the same dimensions as the matrix $\mathbf{D}$. The update formula can then be written as $\bm{\pi}^{(t)\top}=((\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C})\bm{\pi}^{(t-1)\top}$ from which it follows
$\bm{\pi}^{(t)\top}=((\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C})^{t}\bm{\pi}^{(0)\top}.$
Assuming that the dominant eigenvalue of the matrix $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C}$ is simple and has modulus one, the iterate $\bm{\pi}^{(t)}$ converges towards the corresponding eigenvector.
Lemma 4.5.
Consider the update $\overline{\bm{\pi}}^{(t)\top}=((\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{C})\overline{\bm{\pi}}^{(t-1)\top}$ where $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}$ is applied via $k_{t}$ steps of the Richardson iteration method, and $k_{t}$ is monotonically increasing with the number of outer-loop iterations $t$. Then, the sequence $\{\overline{\bm{\pi}}^{(t)\top}\}$ converges to the same limit as the sequence $\{\bm{\pi}^{(t)\top}\}$.
We now establish our main result for Algorithm 2 by considering the case where the diagonal matrix $\mathbf{D}^{(t-1)}$ depends (non-linearly) on $\bm{\pi}^{(t-1)}$, which leads to couplings in the dependency of $\bm{\pi}^{(t)}$ with respect to $\bm{\pi}^{(t-1)}$. Note that this is exactly the scenario of Algorithm 2.
Theorem 4.6.
Suppose Algorithm 2 with exact linear system solutions provides a sequence $\{\bm{\pi}^{(t)\top}\}$ such that $\|\bm{\pi}^{(t)}-\bm{\pi}\|<\|\bm{\pi}^{(t-1)}-\bm{\pi}\|$, $t>1$. Then, the sequence $\{\overline{\bm{\pi}}^{(t)\top}\}$ produced by replacing the exact linear system solutions in Step $\mathbf{5}$ of Algorithm 2 with $k_{t}$ steps of the Richardson iteration method converges to the same limit as $\{\bm{\pi}^{(t)\top}\}$ provided that $k_{t}$ increases with the number of outer-loop iterations $t$.
Starting with an inexact application of the Richardson iteration method and increasing the number of Richardson iteration steps $k_{t}$ with the number of outer-loop iterations $t$ helps to avoid over-solving in the earlier stages of Step $\mathbf{5}$ in Algorithm 2 where the approximate stationary distribution is far from $\bm{\pi}$. For example, the tolerance up to which we solve the linear system in Step $\mathbf{5}$ follows a geometric criterion such as $\gamma^{t}$ or $\rho/(t+1)^{\theta}$ for some $\gamma,\rho\in(0,1)$ and $\theta>1$. We note here that such tradeoffs are well known in numerical linear algebra and arise often when iterative solvers are used in shift-and-invert algorithms for eigenvalue computations [12].
4.3 Mixed-Precision Computation Performance Analysis
Our general classes of algorithmic approaches exploit a combination of advances in mixed-precision computer architectures and iterative approximate computing methods. Although well studied in machine learning and related applications, we note that such advances in mixed-precision computation have received far less consideration in the context of computing the stationary distribution of Markov processes. Beyond significantly reducing computation times, our use of lower precision computation makes it possible to handle much larger systems of linear equations before hitting the memory bandwidth limitations of today’s highly-efficient advanced processors.
More specifically, the performance tradeoff at the heart of our algorithmic approaches concerns, on one hand, the significant reductions in execution times afforded by mixed-precision computation at the expense of inaccuracies in the results and, on the other hand, the ability to mitigate inaccurate computations, further reduce execution times and guarantee convergence afforded by iterative approximate computing methods. Sections 4.1 and 4.2 address the degree to which our algorithmic approaches mitigate the impact of inaccurate computations, reduce execution times, and ensure convergence. The reductions in execution times provided by mixed-precision computation are also quite significant and vary from one computer processor architecture to the next, with increasing trends over time. Generally speaking, there is typically a factor of 2×2$\times$2 × reduction in computation times with respect to the number of operations per second going from $64$-bit precision to $32$-bit precision, with considerably greater factors of reduction going to $16$-bit precision and to $8$-bit precision including as much as orders of magnitude factors of reduction; see, e.g., [23, 2, 1]. Then, since the cost of any additional iterations incurred by our approaches is often minimal in comparison with the significant reductions from mixed-precision computation in each iteration as shown in the above theoretical results and the empirical results in Section 6, our general classes of algorithmic approaches provide tremendous performance improvements over the most efficient existing methods.
It is well understood that the computational performance of the advanced processor architectures of today can be hindered by memory bandwidth bottlenecks when the problem is sufficiently large relative to the available memory of the processor. In particular, rather than realizing the number of operations per second available from the processor in such cases, the computational performance is significantly reduced and instead dictated by the memory bandwidth of the processor architecture. Therefore, by exploiting reduced precision computation, our general classes of algorithmic approaches further enable us to handle significantly larger systems of linear equations in each iteration at the computational performance afforded by the processor architecture, including the mixed-precision reduction factors above. This is well beyond what is possible with existing methods whose performance would become memory-bandwidth limited for much smaller problem instances.
5 Proofs of Main Results
In this section, we provide the proofs of our main theoretical results presented in Section 4. The presentations of a couple of our proofs are reduced with complete proofs of such cases presented in Appendix A.
5.1 Proof of Lemma 4.1
Proof.
We exclude some of the steps and technical details of our proof due to space limitations, referring the reader to Appendix A.1 for such additional details.
From the specific supposition $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$ together with $10<z\leq\|\bm{\pi}_{i}\|_{1}<$ 1, we obtain
$\|\bm{\pi}^{(t-1)}_{i}\|_{1}=\|\bm{\pi}_{i}\|_{1}+O(\varphi)\qquad\mbox{ and }\qquad\frac{\|\bm{\pi}^{(t-1)}_{i}\|_{1}}{\|\bm{\pi}_{i}\|_{1}}=\frac{\|\bm{\pi}_{i}\|_{1}}{\|\bm{\pi}_{i}\|_{1}}+\frac{O(\varphi)}{\|\bm{\pi}_{i}\|_{1}}=1+O(\varphi),$
from which it follows that
$\frac{\bm{\pi}^{(t-1)}_{i}}{\|\bm{\pi}^{(t-1)}_{i}\|_{1}}=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}+O(\varphi),\qquad i\in[m].$
This together with the definition of $\mathbf{R}^{(t-1)}_{ij}:=({\bm{\pi}^{(t-1)}_{i}}/{\|\bm{\pi}^{(t-1)}_{i}\|_{1}})\mathbf{P}_{ij}\mathbf{1}$ from our Algorithm 1 yields
$\mathbf{R}^{(t-1)}_{ij}=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}\mathbf{P}_{ij}\mathbf{1}+O(\varphi)\mathbf{P}_{ij}\mathbf{1},\qquad\forall i,j\in[m],\;i\neq j.$
Then, from the definition of $\mathbf{R}_{ij}:=({\bm{\pi}_{i}}/{\|\bm{\pi}_{i}\|_{1}})\mathbf{P}_{ij}\mathbf{1}$ in the rightmost equation of (8) and since $\mathbf{P}_{ij}=O(\epsilon)$ from (3) for $i\neq j$, we have
$\mathbf{R}^{(t-1)}_{ij}=\mathbf{R}_{ij}+O(\varphi\epsilon),\qquad\forall i,j\in[m],\;i\neq j,$
which renders the desired result for $\mathbf{R}^{(t-1)}$ as an $O(\varphi\epsilon)$ approximation of $\mathbf{R}$ when $i\neq j$. Given that $\mathbf{R}^{(t-1)}$ and $\mathbf{R}$ are both stochastic matrices, we further conclude
$\mathbf{R}^{(t-1)}_{ii}=1-\sum_{j\in[m]:j\neq i}\mathbf{R}^{(t-1)}_{ij}=1-\sum_{j\in[m]:j\neq i}\mathbf{R}_{ij}+O(\varphi\epsilon)=\mathbf{R}_{ii}+O(\varphi\epsilon),\qquad\forall i\in[m],$
thus completing the proof of $\mathbf{R}^{(t-1)}$ as an $O(\varphi\epsilon)$ approximation of $\mathbf{R}$.
Next, from the middle equation in (11), we have
$(\mathbf{I}-\mathbf{R})=\mathbf{S}\begin{pmatrix}0&0\\ 0&\mathbf{I}-\mathbf{B}\end{pmatrix}\mathbf{S}^{-1}.$ (17)
Given the theorem statements $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}=\mathbf{s}\mathbf{R}$, we take the difference between the former statement and the latter statement on both sides and derive
$(\mathbf{s}^{(t-1)}-\mathbf{s})(\mathbf{I}-\mathbf{R})$ $=\mathbf{s}^{(t-1)}(\mathbf{R}^{(t-1)}-\mathbf{R})$ (18)
$(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}\begin{pmatrix}0&0\\ 0&\mathbf{I}-\mathbf{B}\end{pmatrix}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\mathbf{s}^{(t-1)}(\mathbf{R}^{(t-1)}-\mathbf{R})\mathbf{S},$
where (a) follows upon substituting (17) into (18) and multiplying both sides on the right by $\mathbf{S}$. Given the theorem statement $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ (according to the rightmost equation in (9)) and $\|\mathbf{S}\|=O(1)$ from the rightmost equation in (11), together with $\mathbf{R}^{(t-1)}=\mathbf{R}_{ij}+O(\varphi\epsilon)$ established above, we obtain
$(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}\begin{pmatrix}0&0\\ 0&\mathbf{I}-\mathbf{B}\end{pmatrix}=\mathbf{R}^{(t-1)}-\mathbf{R}=O(\varphi\epsilon).$ (19)
Letting $(\hat{{s}}_{1},\hat{{s}}_{2},\ldots,\hat{{s}}_{m})=(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}$, we have
$\hat{{s}}_{1}=((\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S})_{1}=(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{1}$ $=\|\mathbf{s}^{(t-1)}\|_{1}-\|\mathbf{s}\|_{1}=0,$ (20)
$(\hat{{s}}_{2},\ldots,\hat{{s}}_{m})(\mathbf{I}-\mathbf{B})$ =$\stackrel{{\scriptstyle(a)}}{{=}}O(\varphi\epsilon),$
where (a) follows from (19). Since $(\mathbf{I}-\mathbf{B})$ is nonsingular from (12), we first multiply both sides of the last equation on the right by $(\mathbf{I}-\mathbf{B})^{-1}$ and then take the norm on both sides to conclude
$\|(\hat{{s}}_{2},\ldots,\hat{{s}}_{m})\|\leq O(\varphi\epsilon)\|(\mathbf{I}-\mathbf{B})^{-1}\|.$
From (12), we obtain
$(\hat{{s}}_{2},\ldots,\hat{{s}}_{m})=O(\varphi),$ (21)
and upon combining (20) and (21), we have
$(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}=O(\varphi).$
It therefore follows from the last equation in (11) that $(\mathbf{s}^{(t-1)}-\mathbf{s})=O(\varphi)$, thus completing the proof. ∎
5.2 Proof of Theorem 4.2
Proof.
We exclude some of the steps and technical details of our proof due to space limitations, referring the reader to Appendix A.2 for such additional details.
From (1), we have $\bm{\pi}_{i}=\sum_{j=1}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ and $\bm{\pi}_{i}(\mathbf{I}-\mathbf{P}_{ii})=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$. Similarly, from (5) and (6), we obtain
$\mathbf{v}_{i}\mathbf{P}_{ii}=(1-{g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i}\qquad\mbox{ and }\qquad\mathbf{v}_{i}(\mathbf{I}-\mathbf{P}_{ii})=({g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i},$ (22)
and therefore we conclude
$\bm{\pi}_{i}-\mathbf{v}_{i}$ $=\left(\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}+(-{g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i}\right)(\mathbf{I}-\mathbf{P}_{ii})^{-1}.$ (23)
Given that $\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}$ is a nonsingular matrix of order $n_{i}$ and $\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ is a row vector of order $n_{i}$, there exists an order $n_{i}$ row vector
$\mathbf{c}_{i}=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}^{-1},$ (24)
from which it follows that $\mathbf{c}_{i}=O(\epsilon)$ since the norm of $\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}^{-1}$ is bounded by a constant independent of $\epsilon$ and $\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ is $O(\epsilon)$. Then, upon substitution of (24) into (23), we obtain
$\bm{\pi}_{i}-\mathbf{v}_{i}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\mathbf{c}_{i}\begin{pmatrix}({g}_{i}\epsilon+o(\epsilon))^{-1}&0\\ \mathbf{0}&(\mathbf{I}-\mathbf{T}_{i})^{-1}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}-\mathbf{v}_{i}$
$=\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\mathbf{v}_{i}+({c}_{i_{2}},\ldots,{c}_{i_{n_{i}}})(\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}-\mathbf{v}_{i},$ (25)
where (a) follows upon substituting (6) together with (22). Given that $\mathbf{c}_{i}=O(\epsilon)$, $(\mathbf{I}-\mathbf{T}_{i})^{-1}$ is $O(1)$ from (7) and $\mathbf{V}_{i}$ is also $O(1)$ from above, it follows from (25) that
$\bm{\pi}_{i}-\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\mathbf{v}_{i}$ $=O(\epsilon).$
Let ${b}_{i}={c}_{i_{1}}/({g}_{i}\epsilon+o(\epsilon))$. Since $\mathbf{c}_{i}=O(\epsilon)$, this implies that ${c}_{i_{1}}$ is of order ${h}_{i}\epsilon+o(\epsilon)$ for some ${h}_{i}>0$, and thus we have
${b}_{i}=\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}=\frac{{h}_{i}\epsilon+o(\epsilon)}{{g}_{i}\epsilon+o(\epsilon)}=\frac{{h}_{i}}{{g}_{i}}+o(1)=O(1),$
from which the desired result (16) follows.
Next, let the error in the approximate solution $\bm{\pi}^{(t-1)}$ at iteration $t-1$ of Algorithm 1 be of order $\varphi$, namely $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$. Then, from Lemma 4.1, we have
$\mathbf{R}^{(t-1)}=\mathbf{R}+O(\varphi\epsilon)\qquad\mbox{and}\qquad\mathbf{s}^{(t-1)}=\mathbf{s}+O(\varphi).$ (26)
Starting with the $m$-th component of $\bm{\pi}^{(t)}$ for iteration $t$ of Algorithm 1, we derive
$\bm{\pi}_{m}^{(t)}$ $=\bm{\pi}_{m}^{(t)}\mathbf{P}_{mm}+\sum_{j=1}^{m-1}s_{j}^{(t-1)}\hat{\bm{\pi}}_{j}^{(t-1)}\mathbf{P}_{jm}$
$\bm{\pi}_{m}^{(t)}(\mathbf{I}-\mathbf{P}_{mm})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{m-1}(s_{j}+O(\varphi))(\hat{\bm{\pi}}_{j}+O(\varphi))\mathbf{P}_{jm}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\bm{\pi}_{m}(\mathbf{I}-\mathbf{P}_{mm})+\sum_{j=1}^{m-1}O(\varphi)\mathbf{P}_{jm},$ (27)
where (a) follows from substituting the rightmost equation in (26) and from the $O(\varphi)$ approximation of $\bm{\pi}^{(t-1)}$, and (b) follows from (8) and (10) and from $s_{j}$ and $\hat{\bm{\pi}}_{j}$ both being $O(1)$ together with $\varphi<1$. Similarly to (24), given that $\begin{pmatrix}\mathbf{v}_{m}\\ \mathbf{V}_{m}\end{pmatrix}$ is a nonsingular matrix of order $n_{m}$ whose norm is bounded by a constant independent of $\epsilon$, and given that $\sum_{j=1}^{m-1}O(\varphi)\mathbf{P}_{jm}$ is a row vector of order $n_{m}$ whose components are $O(\varphi\epsilon)$, there exists a row vector $\bar{\mathbf{c}}_{m}$ of order $n_{m}$ whose components are $O(\varphi\epsilon)$ such that
$\sum_{j=1}^{m-1}O(\varphi)\mathbf{P}_{jm}=\bar{\mathbf{c}}_{m}\begin{pmatrix}\mathbf{v}_{m}\\ \mathbf{V}_{m}\end{pmatrix}.$ (28)
Multiplying both sides of (27) by $(\mathbf{I}-\mathbf{P}_{mm})^{-1}$ and substituting (28), we obtain
$\bm{\pi}_{m}^{(t)}-\bm{\pi}_{m}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\bar{\mathbf{c}}_{m}\begin{pmatrix}\displaystyle\frac{\mathbf{v}_{m}}{{g}_{m}\epsilon+o(\epsilon)}\vspace*{0.15cm}\\ (\mathbf{I}-\mathbf{T}_{m})^{-1}\mathbf{V}_{m}\end{pmatrix}$
$=\left(\frac{\bar{{c}}_{m_{1}}}{{g}_{m}\epsilon+o(\epsilon)}\right)\mathbf{v}_{m}+(\bar{{c}}_{m_{2}},\ldots,\bar{{c}}_{m_{n_{m}}})(\mathbf{I}-\mathbf{T}_{m})^{-1}\mathbf{V}_{m},$ (29)
where (a) follows upon substitution of (6). From (16), we have
$\mathbf{v}_{m}={b}_{m}^{-1}\bm{\pi}_{m}+O(\epsilon),$ (30)
which together with (29), (30), $\bar{{c}}_{m_{k}}=O(\varphi\epsilon)$, ${c}_{m_{1}}=O(\epsilon)$, $(\mathbf{I}-\mathbf{T}_{m})^{-1}=O(1)$ and $\mathbf{V}_{m}=O(1)$ yields
$\bm{\pi}_{m}^{(t)}-\bm{\pi}_{m}$ $=\left(\frac{\bar{{c}}_{m_{1}}}{{g}_{m}\epsilon+o(\epsilon)}\right)\left(\frac{{g}_{m}\epsilon+o(\epsilon)}{{c}_{m_{1}}}\bm{\pi}_{m}+O(\epsilon)\right)+(\bar{{c}}_{m_{2}},\ldots,\bar{{c}}_{m_{n_{m}}})(\mathbf{I}-\mathbf{T}_{m})^{-1}\mathbf{V}_{m}$
$\bm{\pi}_{m}^{(t)}$ $=(1+O(\varphi))\bm{\pi}_{m}+O(\varphi\epsilon).$
It then follows that
$\hat{\bm{\pi}}_{m}^{(t)}=\frac{\bm{\pi}_{m}^{(t)}}{\|\bm{\pi}_{m}^{(t)}\|_{1}}=\frac{(1+O(\varphi))\bm{\pi}_{m}+O(\varphi\epsilon)}{\|(1+O(\varphi))\bm{\pi}_{m}+O(\varphi\epsilon)\|_{1}}=\frac{\bm{\pi}_{m}}{\|\bm{\pi}_{m}\|_{1}}+O(\varphi\epsilon),$
from which we conclude
$\hat{\bm{\pi}}_{m}^{(t)}=\hat{\bm{\pi}}_{m}+O(\varphi\epsilon).$ (31)
Now, arguing by induction with respect to the base case (31), we can therefore establish
$\frac{\bm{\pi}_{i}^{(t)}}{\|\bm{\pi}_{i}^{(t)}\|_{1}}$ $=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}+O(\varphi\epsilon),\qquad\forall i\in[m].$
It then follows from the initial arguments in the proof of Lemma 4.1 that
$\bm{\pi}^{(t)}=\bm{\pi}+O(\varphi\epsilon),\qquad\forall i\in[m],$
which yields the desired result. ∎
5.3 Proof of Lemma 4.3
Proof.
Under the process employed for selecting the precision of computations in Step $\mathbf{5}$ and Step $\mathbf{3}$ of Algorithm 1 with respect to the condition number and the norm of the matrix $\mathbf{A}$ of the linear system, we know that the assumptions of the mixed-precision analysis of iterative refinement by Moler [20] are satisfied. It then follows from the convergence results of Moler [20] that, under these conditions, the iterative refinement method is guaranteed to converge linearly with an approximation error that decreases by a factor of $O(\kappa(\mathbf{A}))$ in each iteration, thus completing the proof. ∎
5.4 Proof of Lemma 4.4
Proof.
First, notice that the matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ is block-lower triangular, with the $i$-th on-diagonal block being equal to $\mathbf{I}-\mathbf{P}_{ii}$. Since $\mathbf{P}_{ii}$ has only one eigenvalue that is close to one, and since this eigenvalue $\lambda_{i_{1}}$ is equal to $1-{g}_{i}\epsilon+o(\epsilon)$, the smallest eigenvalue of $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ is of the order $O(\epsilon)$. Expressing
$\mathbf{D}^{\top}-\mathbf{U}^{\top}=\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}^{-\top}\mathbf{U}^{\top}),$
and inverting both sides yields
$(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}=(\mathbf{I}-\mathbf{D}^{-\top}\mathbf{U}^{\top})^{-1}\mathbf{D}^{-\top}.$
Assuming the invertibility of $\mathbf{D}^{\top}-\mathbf{U}^{\top}$, it follows that $\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|<1$, and thus $(\mathbf{I}-\mathbf{D}^{-\top}\mathbf{U}^{\top})^{-1}$ possesses a convergent Neumann series that can be expanded as
$(\mathbf{I}-\mathbf{D}^{-\top}\mathbf{U}^{\top})^{-1}=\mathbf{I}+\mathbf{D}^{-\top}\mathbf{U}^{\top}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{2}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{3}+\cdots,$
which leads to
$(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}=\left(\mathbf{I}+\mathbf{D}^{-\top}\mathbf{U}^{\top}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{2}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{3}+\cdots\right)\mathbf{D}^{-\top}.$
Now, upon truncating the above expression for $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}$ up to $k$ terms, we obtain
$(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k]}^{-1}=\left(\mathbf{I}+\mathbf{D}^{-\top}\mathbf{U}^{\top}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{2}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{3}+\cdots+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k}\right)\mathbf{D}^{-\top},$
which results in an error of
$(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}-(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k]}^{-1}=\left((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k+2}+\cdots\right)\mathbf{D}^{-\top}.$
Recalling from above that $\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|<1$, we obtain the inequality
$\|(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}-(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k]}^{-1}\|\leq\|\mathbf{D}^{-\top}\|\dfrac{\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|^{k+1}}{1-\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|}.$
From this upper bound on the error we deduce that, when $\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|\ll 1$, a small value of $k$ can lead to an accurate approximation of the inverse $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}$, which yields the desired result. ∎
5.5 Proof of Lemma 4.5
Proof.
The expression for the update of the lemma can be equivalently written as
$\overline{\bm{\pi}}^{(t)\top}=\prod_{t=1}\left(\left(\left(\mathbf{D}^{\top}-\mathbf{U}^{\top}\right)^{-1}-\left((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots\right)\mathbf{D}^{-\top}\right)\mathbf{L}^{\top}\mathbf{C}\right)\bm{\pi}^{(0)\top}.$
Then, in order for the sequence $\{\overline{\bm{\pi}}^{(t)\top}\}$ to converge to the same limit as the sequence $\{\bm{\pi}^{(t)\top}\}$, we need to have $\prod_{t=1}(((\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}-((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots)\mathbf{D}^{-\top})\mathbf{L}^{\top}\mathbf{C})$ converge to the dominant eigendirection of $(\mathbf{D}^{\top}-\mathbf{U}^{\top})^{-1}\mathbf{L}^{\top}\mathbf{C}$, which is true provided that $k_{t}$ increases monotonically with the number of outer-loop iterations $t$, thus completing the proof. ∎
5.6 Proof of Theorem 4.6
Proof.
Similarly to Lemma 4.5, the fact that $k_{t}$ increases with the number of outer-loop iterations $t$ implies that the errors introduced by the Richardson iteration method decrease with increasing $t$. However, in contrast with Lemma 4.5 where $\mathbf{D}^{(t-1)}$ is fixed, the dependence of $\mathbf{D}^{(t-1)}$ on $\bm{\pi}^{(t-1)}$ implies that the right-hand side of the linear system solutions in Step $\mathbf{5}$ of Algorithm 2 is also perturbed. Nonetheless, the operations involved in $\mathbf{D}^{(t-1)}$ are linear and thus the perturbation introduced at iteration $t$ are of the same order as the error introduced by the Richardson iteration method in the previous iteration $t-1$. Hence, as long as the application of the Richardson iteration method becomes more accurate over time, Algorithm 2 converges to the true stationary distribution $\bm{\pi}$ analogous to the arguments establishing Theorem 4.2.
More specifically, the update at the outer-loop iteration $t$ satisfies
$\overline{\bm{\pi}}^{(t)\top}=(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top}+(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\overline{\mathbf{D}}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top},$
where $\overline{\mathbf{D}}^{(t-1)}$ denotes the part of $\mathbf{D}^{(t-1)}$ attributed to the non-exact solution of the linear system in the previous iteration $t-1$ and is of the order $O\left(\|\mathbf{D}^{-\top}\|\dfrac{\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|^{k_{t-1}+1}}{1-\|\mathbf{D}^{-\top}\mathbf{U}^{\top}\|}\right)$. Ignoring the second term which depends quadratically on the error tolerance used in the Richardson iteration method, we are left with
$\overline{\bm{\pi}}^{(t)\top}$ $\approx(\mathbf{D}^{\top}-\mathbf{U}^{\top})_{[k_{t}]}^{-1}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top}$
$=\left(\mathbf{D}^{\top}-\mathbf{U}^{\top}\right)^{-1}-\left((\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+1}+(\mathbf{D}^{-\top}\mathbf{U}^{\top})^{k_{t}+2}+\cdots\right)\mathbf{D}^{-\top}\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\overline{\bm{\pi}}^{(t-1)\top},$
which eventually reduces to the iterative behavior of Algorithm 1 and thus the convergence follows from the arguments establishing Theorem 4.2. This yields the desired result. ∎
6 Numerical Experiments
In this section we present a representative sample of numerical experiments that support our theoretical results and empirically evaluate our general classes of algorithmic approaches in comparison with the most efficient existing numerical methods for solving NCD Markov processes. Since the most efficient existing algorithms are the iterative numerical methods due to Takahashi [27], Vantilborgh [28], and Koury, McAllister and Stewart [18], with the latter being the most recent of the three methods, and since all three numerical methods exploit aggregation-disaggregation in a very similar manner with similar convergence behaviors [7], we selected the Koury-McAllister-Stewart (KMS) algorithm as the baseline method for the numerical evaluation comparison of our algorithms.
We first present details on the set up of our numerical experiments and then we present and discuss the performance results from a representative sample of our numerical experiments. Beyond quantitatively supporting our theoretical results, these empirical performance results demonstrate that our general classes of algorithmic approaches exhibit relatively little or no increase in the number of outer-loop iterations and orders of magnitude improvements in the computation time, both in comparison with the baseline method of the original KMS algorithm.
6.1 Experimental Setup
Our numerical experiments of solving NCD Markov processes with Algorithm 1, Algorithm 2 and the baseline original KMS algorithm are conducted in Python on a standard compute infrastructure with a $2.5$Ghz i7777 processor and $100$GB of RAM. The $n\times n$ transition probability matrix $\mathbf{P}$ of block structural form (2) for each independent trial run was randomly generated with each element independently and identically sampled from a uniform distribution on the unit interval. Then, to satisfy the conditions (3), we reduce the magnitude of the elements of the off-diagonal block matrices by proportionally scaling these elements such $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m]$, $i\neq j$; and, to satisfy the requirement that $\mathbf{P}$ be a stochastic matrix, we correspondingly and proportionally scale each row of the matrices $\mathbf{P}_{ii}$, $i\in[m]$, so that each row of the resulting matrix $\mathbf{P}$ sum to one.
Next, we execute every step of each of the three algorithms on the above computational infrastructure, computing the number of outer-loop iterations and computing the number of operations for each step, with the latter converted to obtain the computation time. For every algorithm, the number of operations required to compute the solution of a $k\times k$ linear system using LU decomposition is given by $3k^{3}$, with the operations per second depending upon the precision used for the computation. Other than Step $\mathbf{5}$ and Step $\mathbf{3}$ which are performed in lower precision in Algorithm 1 and Algorithm 2 (see below), all three algorithms execute the same number of operations for each outer-loop iteration which are performed in full-precision. The computation time for each aspect of each algorithm is then obtained by multiplying the number of operations by the processor performance for the level of precision used for the computation. For this purpose, we used the performance of the NVIDIA H100 processor [23, 2] whose peak performance for double-, single- and half-precision floating point arithmetic are 34, 67 and 134 Tflops per second, respectively. The baseline original KMS algorithm performs all its computations in double-precision; Algorithm 1 performs the computations of Step $\mathbf{5}$ and Step $\mathbf{3}$ typically in single-precision; Algorithm 2 performs the computations of Step $\mathbf{5}$ in single-precision and the computations of Step $\mathbf{3}$ in full-precision.
The above computation times are only used when the matrix of each linear system fits within the available memory on the processor architecture. In particular, when the size of the matrix $\mathbf{P}_{ii}$ at the level of precision employed exceeds the available memory of $96$GB, then the computation of the linear system becomes memory bandwidth limited as described in Section 4.3. Based on the specification for the NVIDIA H100 processor [23, 2], the memory bandwidth of $3.35$ Tbytes per second is instead used when any of the algorithms exceed the available memory and become bandwidth limited. As noted in Section 4.3, the use of lower precision computation in our general algorithmic approaches makes it possible to handle larger systems of linear equations before hitting the memory bandwidth limitations of the processor architecture.
Lastly, it is important to note that the performance benefits of Algorithm 1 and Algorithm 2 over the baseline original KMS algorithm stem from a combination of the higher performance throughput of the lower precision arithmetic and the reduced operations of the iterative approximate computing methods.
6.2 Empirical Performance Results
We now present the performance results from our numerical experiments of solving NCD Markov processes with Algorithm 1 and Algorithm 2 that support our theoretical results and quantify the empirical benefits of our general algorithmic approaches over the baseline original KMS algorithm. The primary parameters that influence the computation of the solution of (1) to obtain the stationary distribution $\bm{\pi}$ of NCD Markov processes concern: ($a$) the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$; ($b$) the number $m$ of diagonal block matrices; and ($c$) the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$. We consider collections of performance result comparisons from numerical experiments in which two of the primary parameters are fixed and the remaining primary parameter varies over a range of values. Each set of performance result comparisons comprises evaluation of our conservatively modified KMS approach — Algorithm 1 — denoted by KMS+IR, our aggressively modified KMS approach — Algorithm 2 — denoted by KMS+Ri, and the baseline KMS algorithm. The quantitative metrics of interest in each collection of performance comparisons are the mean number of iterations and the mean computation time in seconds, taken over ten independent experimental trial runs.
For Algorithm 1, we select the computational precision level with respect to the condition number and the norm of the matrix of the linear systems of Step $\mathbf{5}$ and Step $\mathbf{3}$ as defined in Section 3.1. We then employ the iterative refinement method as described in Section 3.1 to compute the solution of each linear system to full-precision accuracy. For Algorithm 2, we employ the Richardson iteration method with block-Jacobi preconditioning as described in Section 3.2 to solve the linear systems of Step $\mathbf{5}$ with $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ as the iteration matrix and $\mathbf{L}^{\top}\mathbf{D}^{(t-1)}\bm{\pi}^{(t-1)\top}$ as the right-hand side. In particular, the linear systems of Step $\mathbf{5}$ with the coefficient matrix $\mathbf{D}^{\top}-\mathbf{U}^{\top}$ are preconditioned by the block-diagonal matrix $\mathbf{D}^{\top}$, where the preconditioner is applied via LU factorization computed in $32$-bit precision. For the solution of each linear system of Step $\mathbf{5}$, we start with a number of maximum-allowed iterations equal to ten and increase this value by five per outer iteration. The linear system of Step $\mathbf{3}$ is solved exactly in full precision so that we can focus on the performance benefits of our algorithmic approach for the larger linear systems of Step $\mathbf{5}$.
We present the results from a representative sample of our numerical experiments. Figure 1 presents the set of performance result comparisons for the case where the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are fixed to be $200$, the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is fixed to be $0.1$, and the number $m$ of diagonal block matrices is varied such that $m\in\{5,10,20,50,100\}$. We first observe from Figure 1(a) that the mean number of iterations under KMS+IR are either identical to that of the original KMS algorithm or differ by at most one, with the mean number of iterations under KMS+Ri only somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$, remains fairly flat for all algorithms, which is as expected since the linear increase is in the number of linear systems of the same size being solved. From Figure 1(b), we observe that both of our algorithmic approaches provide orders of magnitude reduction in the mean computation time and that such performance improvements increase with the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$. The mean computation times for each algorithm increase with $m$ as expected due to the linear increase in the number of linear systems of the same size being solved, with the rate of increase greatest for the original KMS algorithm, with much smaller rates of increase for both of our general algorithmic approaches, and with even smaller rates of increase for KMS+Ri than for KMS+IR. We note that our choice for the number of maximum-allowed iterations in the solution of each linear system of Step $\mathbf{5}$ under KMS+Ri leads to the lowest mean computation times in Figure 1(b). The increase in the number $m$ of block matrices $\mathbf{P}_{ii}$, $i\in[m]$, under KMS+Ri increases the part of the iteration matrix not captured by the block-Jacobi preconditioner, but this use of the method becomes computationally less expensive and thus there is an interesting algorithmic tradeoff which can be exploited within the context of specific applications. Meanwhile, the rate of increase under KMS+Ri as a function of the number of block matrices is the smallest among all algorithms.
(a) Mean number of iterations.
(b) Mean computation times (seconds).
Figure 1: Performance comparison of our modified KMS algorithms (KMS+IR, KMS+Ri) and the KMS algorithm for diagonal matrices $\mathbf{P}_{ii}$ with dimension $n_{i}=200$ and off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$ with $\epsilon=0.1$, as a function of the number of blocks $m$. Performance results represent averages over $10$ independent trial runs.
Figure 2 presents the set of performance result comparisons for the case where the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is fixed to be $0.1$, and the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are varied such that $n_{i}\in\{100,200,500,1000\}$. We first observe from Figure 2(a) that, once again, the mean number of iterations under KMS+IR are essentially identical to that of the original KMS algorithm, with the mean number of iterations under KMS+Ri only somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, remains relatively flat for all algorithms, which is as expected due to the shared properties of the outer loop of all three algorithms. From Figure 2(b), we observe that both of our algorithmic approaches provide orders of magnitude reduction in the mean computation time and that such performance improvements are consistent and increasing across the various dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$. The mean computation times for each algorithm increase with $n_{i}$ as expected due to the linear increase in the size of the linear systems being solved, with the rate of increase somewhat larger for the original KMS algorithm, with somewhat smaller rates of increase for both of our general algorithmic approaches, and with somewhat smaller rates of increase for KMS+Ri than for KMS+IR. Once again, we note that our general KMS+Ri approach provides the lowest mean computation times for the reasons described above, although the differences with our general KMS+IR approach is relatively small but grows somewhat with increases in the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$.
(a) Mean number of iterations.
(b) Mean computation times (seconds).
Figure 2: Performance comparison of our modified KMS algorithms (KMS+IR, KMS+Ri) and the KMS algorithm for number of blocks $m=20$ and off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$ with $\epsilon=0.1$, as a function of the dimension $n_{i}$ of the diagonal matrices $\mathbf{P}_{ii}$. Performance results represent averages over $10$ independent runs.
Figure 3 presents the set of performance result comparisons for the case where the dimensions $n_{i}$ of the diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, are fixed to be $200$, the number $m$ of diagonal block matrices $\mathbf{P}_{ii}$, $i\in[m]$, is fixed to be $20$, and the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, is varied such that $\epsilon\in\{0.01,0.05,0.1,0.15,0.2\}$. We first observe from Figure 3(a) that the mean number of iterations under KMS+IR are identical to that of the original KMS algorithm, with the mean number of iterations under KMS+Ri somewhat higher requiring a few additional iterations (though of lower computational complexity). The mean number of iterations as a function of the magnitude $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$, remains relatively flat for both KMS and KMS+IR, as expected again due to the shared properties of the outer loop of all three algorithms. However, as observed in Figure 3(a), the mean number of iterations under KMS+Ri increases and requires more outer-loop iterations as $\epsilon$ increases. Due to the nature of the block-Jacobi preconditioner, the convergence of the Richardson iteration method becomes slower with increasing $\epsilon$ because now the part of the iteration matrix that is left outside the block-diagonal structure becomes more computationally significant. From Figure 3(b), we observe that both of our algorithmic approaches provide orders of magnitude reduction in the mean computation time with such performance improvements remaining flat across the various magnitudes $\epsilon$ of transitions in the off-diagonal block matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$, $i,j\in[m],i\neq j$. The mean computation times for each algorithm remain relatively flat as a function of the magnitude $\epsilon$, which is as expected due to the shared properties of the outer loop of all three algorithms. We note that our general KMS+Ri approach provides somewhat lower mean computations times than KMS+IR for smaller values of $\epsilon$ and that KMS+IR provides somewhat lower mean computations times than KMS+Ri for larger values of $\epsilon$. This is due to the nature of the block-Jacobi preconditioner, as described above.
(a) Mean number of iterations.
(b) Mean computation times (seconds).
Figure 3: Performance comparison of our modified KMS algorithms (KMS+IR, KMS+Ri) and the KMS algorithm for diagonal matrices $\mathbf{P}_{ii}$ with dimension $n_{i}=200$ and number of blocks $m=20$, as a function of $\epsilon$ for the off-diagonal matrices $\|\mathbf{P}_{ij}\|=O(\epsilon)$. Performance results represent averages over $10$ independent trial runs.
7 Conclusions
We consider in this paper the problem of computing the stationary distribution of NCD Markov processes, a well-established area in the classical theory of Markov processes with broad applications in the design, modeling, analysis and optimization of computer systems. Our focus is on the design and analysis of numerical methods that address the performance bottlenecks of existing approaches and provide significant computational improvements to support the very large scale of current and emerging computer systems and applications as well as their potential deployment in online or real-time environments. In particular, we design general classes of algorithmic solution approaches that exploit forms of mixed-precision computation to significantly reduce computation times and exploit forms of iterative approximate methods to address the impact of inaccurate computations, further reduce computation times, and guarantee convergence. We then derive a mathematical analysis of our general classes of algorithmic approaches that rigorously establishes theoretical results on approximation errors, convergence behaviors, and other algorithmic properties. Numerical experiments empirically demonstrate that our general algorithmic approaches provide orders of magnitude improvements in the computation times to determine the stationary distribution of NCD Markov processes over the most efficient existing numerical methods which are based on aggregation and disaggregation.
Appendix A Full Proofs of Some Theoretical Results
We present in this appendix complete proofs of some of the theoretical results whose corresponding proofs in Section 5 excluded some technical details and steps.
A.1 Proof of Lemma 4.1
Proof.
From the specific supposition $\bm{\pi}^{(t-1)}_{i}=\bm{\pi}_{i}+O(\varphi)$ together with $10<z\leq\|\bm{\pi}_{i}\|_{1}<$ 1, we obtain
$\|\bm{\pi}^{(t-1)}_{i}\|_{1}=\|\bm{\pi}_{i}\|_{1}+O(\varphi)\qquad\mbox{ and }\qquad\frac{\|\bm{\pi}^{(t-1)}_{i}\|_{1}}{\|\bm{\pi}_{i}\|_{1}}=\frac{\|\bm{\pi}_{i}\|_{1}}{\|\bm{\pi}_{i}\|_{1}}+\frac{O(\varphi)}{\|\bm{\pi}_{i}\|_{1}}=1+O(\varphi),$
from which it follows that
$\frac{\bm{\pi}^{(t-1)}_{i}}{\|\bm{\pi}^{(t-1)}_{i}\|_{1}}=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}+O(\varphi),\qquad i\in[m].$
This together with the definition of $\mathbf{R}^{(t-1)}_{ij}:=({\bm{\pi}^{(t-1)}_{i}}/{\|\bm{\pi}^{(t-1)}_{i}\|_{1}})\mathbf{P}_{ij}\mathbf{1}$ from our Algorithm 1 yields
$\mathbf{R}^{(t-1)}_{ij}=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}\mathbf{P}_{ij}\mathbf{1}+O(\varphi)\mathbf{P}_{ij}\mathbf{1},\qquad\forall i,j\in[m],\;i\neq j.$
Then, from the definition of $\mathbf{R}_{ij}:=({\bm{\pi}_{i}}/{\|\bm{\pi}_{i}\|_{1}})\mathbf{P}_{ij}\mathbf{1}$ in the rightmost equation of (8) and since $\mathbf{P}_{ij}=O(\epsilon)$ from (3) for $i\neq j$, we have
$\mathbf{R}^{(t-1)}_{ij}=\mathbf{R}_{ij}+O(\varphi\epsilon),\qquad\forall i,j\in[m],\;i\neq j,$
which renders the desired result for $\mathbf{R}^{(t-1)}$ as an $O(\varphi\epsilon)$ approximation of $\mathbf{R}$ when $i\neq j$. Given that $\mathbf{R}^{(t-1)}$ and $\mathbf{R}$ are both stochastic matrices, we further conclude
$\mathbf{R}^{(t-1)}_{ii}=1-\sum_{j\in[m]:j\neq i}\mathbf{R}^{(t-1)}_{ij}=1-\sum_{j\in[m]:j\neq i}\mathbf{R}_{ij}+O(\varphi\epsilon)=\mathbf{R}_{ii}+O(\varphi\epsilon),\qquad\forall i\in[m],$
thus completing the proof of $\mathbf{R}^{(t-1)}$ as an $O(\varphi\epsilon)$ approximation of $\mathbf{R}$.
Next, from the middle equation in (11), we have
$(\mathbf{I}-\mathbf{R})=\mathbf{S}\begin{pmatrix}0&0\\ 0&\mathbf{I}-\mathbf{B}\end{pmatrix}\mathbf{S}^{-1}.$ (17)
Given the theorem statements $\mathbf{s}^{(t-1)}=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}$ and $\mathbf{s}=\mathbf{s}\mathbf{R}$, we take the difference between the former statement and the latter statement on both sides and derive
$\mathbf{s}^{(t-1)}-\mathbf{s}$ $=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}-\mathbf{s}\mathbf{R}$
$\mathbf{s}^{(t-1)}-\mathbf{s}$ $=\mathbf{s}^{(t-1)}\mathbf{R}^{(t-1)}-\mathbf{s}^{(t-1)}\mathbf{R}+\mathbf{s}^{(t-1)}\mathbf{R}-\mathbf{s}\mathbf{R}$
$(\mathbf{s}^{(t-1)}-\mathbf{s})(\mathbf{I}-\mathbf{R})$ $=\mathbf{s}^{(t-1)}(\mathbf{R}^{(t-1)}-\mathbf{R})$ (18)
$(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}\begin{pmatrix}0&0\\ 0&\mathbf{I}-\mathbf{B}\end{pmatrix}$ =$\stackrel{{\scriptstyle(a)}}{{=}}\mathbf{s}^{(t-1)}(\mathbf{R}^{(t-1)}-\mathbf{R})\mathbf{S},$
where (a) follows upon substituting (17) into (18) and multiplying both sides on the right by $\mathbf{S}$. Given the theorem statement $\|\mathbf{s}\|_{1}=\|\mathbf{s}^{(t-1)}\|_{1}=1$ (according to the rightmost equation in (9)) and $\|\mathbf{S}\|=O(1)$ from the rightmost equation in (11), together with $\mathbf{R}^{(t-1)}=\mathbf{R}_{ij}+O(\varphi\epsilon)$ established above, we obtain
$(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}\begin{pmatrix}0&0\\ 0&\mathbf{I}-\mathbf{B}\end{pmatrix}=\mathbf{R}^{(t-1)}-\mathbf{R}=O(\varphi\epsilon).$ (19)
Letting $(\hat{{s}}_{1},\hat{{s}}_{2},\ldots,\hat{{s}}_{m})=(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}$, we have
$\hat{{s}}_{1}=((\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S})_{1}=(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{1}$ $=\|\mathbf{s}^{(t-1)}\|_{1}-\|\mathbf{s}\|_{1}=0,$ (20)
$(\hat{{s}}_{2},\ldots,\hat{{s}}_{m})(\mathbf{I}-\mathbf{B})$ =$\stackrel{{\scriptstyle(a)}}{{=}}O(\varphi\epsilon),$
where (a) follows from (19). Since $(\mathbf{I}-\mathbf{B})$ is nonsingular from (12), we first multiply both sides of the last equation on the right by $(\mathbf{I}-\mathbf{B})^{-1}$ and then take the norm on both sides to conclude
$\|(\hat{{s}}_{2},\ldots,\hat{{s}}_{m})\|\leq O(\varphi\epsilon)\|(\mathbf{I}-\mathbf{B})^{-1}\|.$
From (12), we obtain
$(\hat{{s}}_{2},\ldots,\hat{{s}}_{m})=O(\varphi),$ (21)
and upon combining (20) and (21), we have
$(\mathbf{s}^{(t-1)}-\mathbf{s})\mathbf{S}=O(\varphi).$
It therefore follows from the last equation in (11) that $(\mathbf{s}^{(t-1)}-\mathbf{s})=O(\varphi)$, thus completing the proof. ∎
A.2 Proof of Theorem 4.2
Proof.
From (1), we have
$\bm{\pi}_{i}=\sum_{j=1}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}\qquad\mbox{ and }\qquad\bm{\pi}_{i}(\mathbf{I}-\mathbf{P}_{ii})=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}.$
Similarly, from (5) and (6), we obtain
$\mathbf{v}_{i}\mathbf{P}_{ii}=(1-{g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i}\qquad\mbox{ and }\qquad\mathbf{v}_{i}(\mathbf{I}-\mathbf{P}_{ii})=({g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i},$ (22)
and therefore we conclude
$(\bm{\pi}_{i}-\mathbf{v}_{i})(\mathbf{I}-\mathbf{P}_{ii})$ $=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}+(-{g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i}$
$\bm{\pi}_{i}-\mathbf{v}_{i}$ $=\left(\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}+(-{g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i}\right)(\mathbf{I}-\mathbf{P}_{ii})^{-1}.$ (23)
Given that $\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}$ is a nonsingular matrix of order $n_{i}$ and $\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ is a row vector of order $n_{i}$, there exists an order $n_{i}$ row vector
$\mathbf{c}_{i}=\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}^{-1},$ (24)
from which it follows that $\mathbf{c}_{i}=O(\epsilon)$ since the norm of $\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}^{-1}$ is bounded by a constant independent of $\epsilon$ and $\sum_{j=1:j\neq i}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}$ is $O(\epsilon)$. Then, upon substitution of (24) into (23), we obtain
$\bm{\pi}_{i}-\mathbf{v}_{i}$ $=\mathbf{c}_{i}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})^{-1}+(-{g}_{i}\epsilon+o(\epsilon))\mathbf{v}_{i}(\mathbf{I}-\mathbf{P}_{ii})^{-1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\mathbf{c}_{i}\begin{pmatrix}({g}_{i}\epsilon+o(\epsilon))^{-1}&0\\ \mathbf{0}&(\mathbf{I}-\mathbf{T}_{i})^{-1}\end{pmatrix}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}-\mathbf{v}_{i}$
$=\mathbf{c}_{i}\begin{pmatrix}\displaystyle\frac{\mathbf{v}_{i}}{{g}_{i}\epsilon+o(\epsilon)}\vspace*{0.15cm}\\ (\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}\end{pmatrix}-\mathbf{v}_{i}$
$=\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\mathbf{v}_{i}+({c}_{i_{2}},\ldots,{c}_{i_{n_{i}}})(\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}-\mathbf{v}_{i},$ (25)
where (a) follows upon substituting (6) together with (22). Given that $\mathbf{c}_{i}=O(\epsilon)$, $(\mathbf{I}-\mathbf{T}_{i})^{-1}$ is $O(1)$ from (7) and $\mathbf{V}_{i}$ is also $O(1)$ from above, it follows from (25) that
$\bm{\pi}_{i}$ $=\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\mathbf{v}_{i}+({c}_{i_{2}},\ldots,{c}_{i_{n_{i}}})(\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}$
$\bm{\pi}_{i}-\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\mathbf{v}_{i}$ $=({c}_{i_{2}},\ldots,{c}_{i_{n_{i}}})(\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}$
$\bm{\pi}_{i}-\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\mathbf{v}_{i}$ $=O(\epsilon).$
Let ${b}_{i}={c}_{i_{1}}/({g}_{i}\epsilon+o(\epsilon))$. Since $\mathbf{c}_{i}=O(\epsilon)$, this implies that ${c}_{i_{1}}$ is of order ${h}_{i}\epsilon+o(\epsilon)$ for some ${h}_{i}>0$, and thus we have
${b}_{i}=\frac{{c}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}=\frac{{h}_{i}\epsilon+o(\epsilon)}{{g}_{i}\epsilon+o(\epsilon)}=\frac{{h}_{i}}{{g}_{i}}+o(1)=O(1),$
from which the desired result (16) follows.
Next, let the error in the approximate solution $\bm{\pi}^{(t-1)}$ at iteration $t-1$ of Algorithm 1 be of order $\varphi$, namely $\bm{\pi}^{(t-1)}=\bm{\pi}+O(\varphi)$. Then, from Lemma 4.1, we have
$\mathbf{R}^{(t-1)}=\mathbf{R}+O(\varphi\epsilon)\qquad\mbox{and}\qquad\mathbf{s}^{(t-1)}=\mathbf{s}+O(\varphi).$ (26)
Starting with the $m$-th component of $\bm{\pi}^{(t)}$ for iteration $t$ of Algorithm 1, we derive
$\bm{\pi}_{m}^{(t)}$ $=\bm{\pi}_{m}^{(t)}\mathbf{P}_{mm}+\sum_{j=1}^{m-1}s_{j}^{(t-1)}\hat{\bm{\pi}}_{j}^{(t-1)}\mathbf{P}_{jm}$
$\bm{\pi}_{m}^{(t)}(\mathbf{I}-\mathbf{P}_{mm})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{m-1}(s_{j}+O(\varphi))(\hat{\bm{\pi}}_{j}+O(\varphi))\mathbf{P}_{jm}$
$=\sum_{j=1}^{m-1}s_{j}\hat{\bm{\pi}}_{j}\mathbf{P}_{jm}+\sum_{j=1}^{m-1}(s_{j}O(\varphi)+\hat{\bm{\pi}}_{j}O(\varphi)+O(\varphi^{2}))\mathbf{P}_{jm}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\bm{\pi}_{m}(\mathbf{I}-\mathbf{P}_{mm})+\sum_{j=1}^{m-1}O(\varphi)\mathbf{P}_{jm},$ (27)
where (a) follows from substituting the rightmost equation in (26) and from the $O(\varphi)$ approximation of $\bm{\pi}^{(t-1)}$, and (b) follows from (8) and (10) and from $s_{j}$ and $\hat{\bm{\pi}}_{j}$ both being $O(1)$ together with $\varphi<1$. Similarly to (24), given that $\begin{pmatrix}\mathbf{v}_{m}\\ \mathbf{V}_{m}\end{pmatrix}$ is a nonsingular matrix of order $n_{m}$ whose norm is bounded by a constant independent of $\epsilon$, and given that $\sum_{j=1}^{m-1}O(\varphi)\mathbf{P}_{jm}$ is a row vector of order $n_{m}$ whose components are $O(\varphi\epsilon)$, there exists a row vector $\bar{\mathbf{c}}_{m}$ of order $n_{m}$ whose components are $O(\varphi\epsilon)$ such that
$\sum_{j=1}^{m-1}O(\varphi)\mathbf{P}_{jm}=\bar{\mathbf{c}}_{m}\begin{pmatrix}\mathbf{v}_{m}\\ \mathbf{V}_{m}\end{pmatrix}.$ (28)
Multiplying both sides of (27) by $(\mathbf{I}-\mathbf{P}_{mm})^{-1}$ and substituting (28), we obtain
$\bm{\pi}_{m}^{(t)}-\bm{\pi}_{m}$ $=\bar{\mathbf{c}}_{m}\begin{pmatrix}\mathbf{v}_{m}\\ \mathbf{V}_{m}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{mm})^{-1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\bar{\mathbf{c}}_{m}\begin{pmatrix}\displaystyle\frac{\mathbf{v}_{m}}{{g}_{m}\epsilon+o(\epsilon)}\vspace*{0.15cm}\\ (\mathbf{I}-\mathbf{T}_{m})^{-1}\mathbf{V}_{m}\end{pmatrix}$
$=\left(\frac{\bar{{c}}_{m_{1}}}{{g}_{m}\epsilon+o(\epsilon)}\right)\mathbf{v}_{m}+(\bar{{c}}_{m_{2}},\ldots,\bar{{c}}_{m_{n_{m}}})(\mathbf{I}-\mathbf{T}_{m})^{-1}\mathbf{V}_{m},$ (29)
where (a) follows upon substitution of (6). From (16), we have
$\mathbf{v}_{m}={b}_{m}^{-1}\bm{\pi}_{m}+O(\epsilon),$ (30)
which together with (29), (30), $\bar{{c}}_{m_{k}}=O(\varphi\epsilon)$, ${c}_{m_{1}}=O(\epsilon)$, $(\mathbf{I}-\mathbf{T}_{m})^{-1}=O(1)$ and $\mathbf{V}_{m}=O(1)$ yields
$\bm{\pi}_{m}^{(t)}-\bm{\pi}_{m}$ $=\left(\frac{\bar{{c}}_{m_{1}}}{{g}_{m}\epsilon+o(\epsilon)}\right)\left(\frac{{g}_{m}\epsilon+o(\epsilon)}{{c}_{m_{1}}}\bm{\pi}_{m}+O(\epsilon)\right)+(\bar{{c}}_{m_{2}},\ldots,\bar{{c}}_{m_{n_{m}}})(\mathbf{I}-\mathbf{T}_{m})^{-1}\mathbf{V}_{m}$
$=O(\varphi)\bm{\pi}_{m}+\frac{\bar{{c}}_{m_{1}}\cdot O(\epsilon)}{{g}_{m}\epsilon+o(\epsilon)}+O(\varphi\epsilon)$
$\bm{\pi}_{m}^{(t)}$ $=(1+O(\varphi))\bm{\pi}_{m}+O(\varphi\epsilon).$
It then follows that
$\hat{\bm{\pi}}_{m}^{(t)}=\frac{\bm{\pi}_{m}^{(t)}}{\|\bm{\pi}_{m}^{(t)}\|_{1}}=\frac{(1+O(\varphi))\bm{\pi}_{m}+O(\varphi\epsilon)}{\|(1+O(\varphi))\bm{\pi}_{m}+O(\varphi\epsilon)\|_{1}}=\frac{\bm{\pi}_{m}}{\|\bm{\pi}_{m}\|_{1}}+O(\varphi\epsilon),$
from which we conclude
$\hat{\bm{\pi}}_{m}^{(t)}=\hat{\bm{\pi}}_{m}+O(\varphi\epsilon).$ (31)
Now, arguing by induction with respect to the base case (31), suppose
$\hat{\bm{\pi}}_{k}^{(t)}=\hat{\bm{\pi}}_{k}+O(\varphi\epsilon)$ (32)
holds for all $k=m,m-1,\ldots,i+1$. Then, for the $i$-th component of $\bm{\pi}^{(t)}$ for iteration $t$ of Algorithm 1, we derive
$\bm{\pi}_{i}^{(t)}$ $=\bm{\pi}_{i}^{(t)}\mathbf{P}_{ii}+\sum_{j=1}^{i-1}s_{j}^{(t-1)}\hat{\bm{\pi}}_{j}^{(t-1)}\mathbf{P}_{ji}+\sum_{j=i+1}^{m}\bm{\pi}_{j}^{(t)}\mathbf{P}_{ji}$
$\bm{\pi}_{i}^{(t)}(\mathbf{I}-\mathbf{P}_{ii})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sum_{j=1}^{i-1}(s_{j}+O(\varphi))(\hat{\bm{\pi}}_{j}+O(\varphi))\mathbf{P}_{ji}+\sum_{j=i+1}^{m}(\bm{\pi}_{j}+O(\varphi\epsilon))\mathbf{P}_{ji}$
$=\sum_{j=1}^{i-1}s_{j}\hat{\bm{\pi}}_{j}\mathbf{P}_{ji}+\sum_{j=1}^{i-1}(s_{j}O(\varphi)+\hat{\bm{\pi}}_{j}O(\varphi)+O(\varphi^{2}))\mathbf{P}_{ji}+\sum_{j=i+1}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}+\sum_{j=i+1}^{m}O(\varphi\epsilon)\mathbf{P}_{ji}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{j=1}^{i-1}\bm{\pi}_{j}\mathbf{P}_{ji}+\sum_{j=i+1}^{m}\bm{\pi}_{j}\mathbf{P}_{ji}+\sum_{j=1}^{i-1}(s_{j}O(\varphi)+\hat{\bm{\pi}}_{j}O(\varphi)+O(\varphi^{2}))\mathbf{P}_{ji}+\sum_{j=i+1}^{m}O(\varphi\epsilon)\mathbf{P}_{ji}$
=$\stackrel{{\scriptstyle(c)}}{{=}}\bm{\pi}_{i}(\mathbf{I}-\mathbf{P}_{ii})+\sum_{j=1:j\neq i}^{m}O(\varphi)\mathbf{P}_{ji},$ (33)
where: (a) follows from substituting the rightmost equation in (26), from the $O(\varphi)$ approximation of $\bm{\pi}^{(t-1)}$, and from substituting (32); (b) follows from (8) and (10); and (c) follows from $s_{j}$ and $\hat{\bm{\pi}}_{j}$ both being $O(1)$ together with $\epsilon<\varphi<1$. Similarly to (28), given that $\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}$ is a nonsingular matrix of order $n_{i}$ whose norm is bounded by a constant independent of $\epsilon$, and given that $\sum_{j=1:j\neq i}^{m}O(\varphi)\mathbf{P}_{ji}$ is a row vector of order $n_{i}$ whose components are $O(\varphi\epsilon)$, there exists a row vector $\bar{\mathbf{c}}_{i}$ of order $n_{i}$ whose components are $O(\varphi\epsilon)$ such that
$\sum_{j=1:j\neq i}^{m}O(\varphi)\mathbf{P}_{ji}=\bar{\mathbf{c}}_{i}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}.$ (34)
Multiplying both sides of (33) by $(\mathbf{I}-\mathbf{P}_{ii})^{-1}$ and substituting (34), we obtain
$\bm{\pi}_{i}^{(t)}-\bm{\pi}_{i}$ $=\bar{\mathbf{c}}_{i}\begin{pmatrix}\mathbf{v}_{i}\\ \mathbf{V}_{i}\end{pmatrix}(\mathbf{I}-\mathbf{P}_{ii})^{-1}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\bar{\mathbf{c}}_{i}\begin{pmatrix}\displaystyle\frac{\mathbf{v}_{i}}{{g}_{i}\epsilon+o(\epsilon)}\vspace*{0.15cm}\\ (\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}\end{pmatrix}$
$=\left(\frac{\bar{{c}}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\right)\mathbf{v}_{i}+(\bar{{c}}_{i_{2}},\ldots,\bar{{c}}_{i_{n_{i}}})(\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i},$ (35)
where (a) follows upon substitution of (6). From (16), we have
$\mathbf{v}_{i}={b}_{i}^{-1}\bm{\pi}_{i}+O(\epsilon),$ (36)
which together with (35), (36), $\bar{{c}}_{i_{k}}=O(\varphi\epsilon)$, ${c}_{i_{1}}=O(\epsilon)$, $(\mathbf{I}-\mathbf{T}_{i})^{-1}=O(1)$ and $\mathbf{V}_{i}=O(1)$ yields
$\bm{\pi}_{i}^{(t)}-\bm{\pi}_{i}$ $=\left(\frac{\bar{{c}}_{i_{1}}}{{g}_{i}\epsilon+o(\epsilon)}\right)\left(\frac{{g}_{i}\epsilon+o(\epsilon)}{{c}_{i_{1}}}\bm{\pi}_{i}+O(\epsilon)\right)+(\bar{{c}}_{i_{2}},\ldots,\bar{{c}}_{i_{n_{i}}})(\mathbf{I}-\mathbf{T}_{i})^{-1}\mathbf{V}_{i}$
$=O(\varphi)\bm{\pi}_{i}+\frac{\bar{{c}}_{i_{1}}\cdot O(\epsilon)}{{g}_{i}\epsilon+o(\epsilon)}+O(\varphi\epsilon)$
$\bm{\pi}_{i}^{(t)}$ $=(1+O(\varphi))\bm{\pi}_{i}+O(\varphi\epsilon).$
It then follows that
$\hat{\bm{\pi}}_{i}^{(t)}=\frac{\bm{\pi}_{i}^{(t)}}{\|\bm{\pi}_{i}^{(t)}\|_{1}}=\frac{(1+O(\varphi))\bm{\pi}_{i}+O(\varphi\epsilon)}{\|(1+O(\varphi))\bm{\pi}_{i}+O(\varphi\epsilon)\|_{1}}=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}+O(\varphi\epsilon),$
from which we conclude
$\hat{\bm{\pi}}_{i}^{(t)}=\hat{\bm{\pi}}_{i}+O(\varphi\epsilon).$
By induction, we therefore have
$\frac{\bm{\pi}_{i}^{(t)}}{\|\bm{\pi}_{i}^{(t)}\|_{1}}$ $=\frac{\bm{\pi}_{i}}{\|\bm{\pi}_{i}\|_{1}}+O(\varphi\epsilon),\qquad\forall i\in[m].$
It then follows from the initial arguments in the proof of Lemma 4.1 that
$\bm{\pi}^{(t)}=\bm{\pi}+O(\varphi\epsilon),\qquad\forall i\in[m],$
which yields the desired result. ∎